OpenStudy (anonymous):

what is the period of the function sin^2(2x)?

OpenStudy (anonymous):

i know that sin has a period of 2pi

OpenStudy (anonymous):

@Shonengetsu @UnkleRhaukus

OpenStudy (anonymous):

period of sine is \(2\pi\) but period of \(\sin(bx)\) is \(\frac{2\pi}{b}\)

OpenStudy (anonymous):

yes but this is sin^2 not sin

OpenStudy (anonymous):

ooooooooooooh so it is

OpenStudy (anonymous):

sorry period in this case would be \(\frac{\pi}{2}\)

OpenStudy (unklerhaukus):

negatives will be positive so the period of sin^2x is half of sinx

OpenStudy (anonymous):

what @UnkleRhaukus said

OpenStudy (anonymous):

half, and then half again

OpenStudy (anonymous):

oh okay thank you!